A total of eight (8) earthquakes have been located for this period. These were scattered throughout the south of the island, from 0 to 5 km depth. Seismicity overnight was dominated by several long-period earthquakes, probably associated with small explosions within the crater. The small, frequent long-period earthquakes recorded over the last few days continued until at least 08:00 today. Since then, mud flow signals have been recorded continuously at the Gages seismic station, and these signals have masked the small long-period earthquakes. The mudflow signals have varied in amplitude throughout the day, and have been recorded at the Bethel and Long Ground stations also. These mudflow signals are due to the large runoff in the ghauts associated with the heavy rainfall that occurred throughout the day.
No EDM measurements were made today, because of the persistent low cloud and heavy rain. GPS measurements were made at St George's Hill, Reid's Hill and Harris' Lookout. The results from yesterday's measurements showed no change from the measurements made on 16 Dec.
Visual observations were made from the helicopter last night. A greater number of glowing rocks were observed in the crater area than had been seen during previous night flights. A large, glowing crack was observed in the lava dome, and small hot rocks were falling continuously from this crack. Spalling of material revealed fresh glowing material. A large block of this material rolled into the crater moat and disintegrated. These observations indicate that the domes in the crater area are highly active and are probably still growing. At dawn this morning a large steam plume was being emitted from the crater area. Visual observations have not been possible since then.